For your retirement planning, you are currently depositing $350 per month into an account that earns...

  1. For your retirement planning, you are currently depositing $350 per month into an account that earns an 8% return, compounded monthly. In 12 years, you expect to increase that deposit by $250 per month (to a total of $600 per month). You plan to retire in 40 years. After you retire, you will move the money into a safe account that earns a guaranteed 3.5% per year.
    1. How much will you have when you retire?
    2. If you expect to live 35 years past retirement, how much will you be able to withdraw per month to live off of?
    3. If you prefer to live off of the interest, thus creating a perpetuity and never touching the principal, how much will you be able to withdraw per month to live off of?
Answer #1

The answers are as follows:

We can solve this question using Financial Calculator CMPD Set: Begin / End. N= Number of Periods. 1./ - Interest pere periodSolution to ca. Set: Begin n = 12X12 - 144 Il. =8+12 - 0.6667 PV. O PMT=-350 FV = 84739.12154 SOLVE Now, Set: Begin Da (40-12Solution to (6): Since the Ethective gate is 3.5 p.a., we have to Convert it into Nominal Interest rate in ond Ive amount reqSolution to (c): Set: Begin no (Doesnot matter) [Suppose 99812-118 1% = 3.445078463:12 -0.287089872 PV7-1544220.497 PMT: 4420